Experimental Particle Physics

1. Research on a Frontier

2. The Current Paradigm—The Standard Model 3. Types of Experiments

4. The Next Step

Research on a Frontier

Particle Physics is a study of the very smallest structure

• Constituents

• Forces

Field has existed since ∼ 1950 − 1960

• Most of the easy “stuff” has been done!

• Experiments are technically more challenging

Field has made progress by pushing on paradigms

• Use the existing “best” theories to decide what phenomena is most interesting

• How you can learn more about the physics

Standard Model of Elementary Particles

Particle physics before ∼ 1960 had established

• Elementary particles appeared to be the basic constituents

of universe electron (e±) proton (p) neutron (n)

• Four fundamental interactions or forces:

1. Strong Force—bound nucleons (p, n) together

2. Electromagnetism—bound e− to nuclei

3. Weak Force—responsible for neutron β decay

4. Gravity—most obvious, but also the weakest

• Host of other “elementary particles”

– neutrinos (νe, νμ)

– muon (μ)

– lots of other hadrons (π∘, K±, Δ’s,...) Seemed to be spare wheels....

Types of Experiments

There are three “classes” of particle physics experiments

1. Fixed target experiments

2. Colliding beam accelerators

3. Non-accelerator experiments

In each case, one wants to observe a particle “interaction”

• requires the initial state to be prepared

• requires a device to detect the results of interaction

• requires tools to analyze the results

Non-Accelerator Experiments

There are some phenomena that can’t be easily observed in an accelerator environment

• Some Grand-Unified Theories predict protons to decay – required tons of protons

– very quiet environment

⇒ place vat of water deep underground

• Study neutrinos from the sun and supernovae ⇒ Sudbury

Neutrino Observatory

Future of Experimental Particle Physics

The next generation of colliders will probe even higher energies

1. SSC in Texas

2. LHC in Geneva, SW

Studies of detailed phenomena

• Rare decays of K and B mesons

• Search for new (and rare) particles

• Understand spectroscopy